# Learning to Generate Unit Tests for Automated Debugging

### v2

@@ 3.3 Generated Unit Tests for Debugging
Building on past work that has shown the benefits of scaling up inference-time compute for LLMs (Wang et al., 2022; Lightman et al., 2023; Snell et al., 2024), we improve UT output accuracy by allocating additional computation to the problem. Specifically, we use self-consistency (SC; Wang et al., 2022), whereby, for a given UT input, we sample output completions (including CoT rationales) and take the most common final UT output (majority vote) as the final answer, as shown in Fig. 3 (top-right). To further boost output accuracy, consistent with Prasad et al. (2024), we upsample UT inputs and only retain those where the final answer gets over 50% of the votes (i.e., 4 votes), discarding the unit test otherwise.
@@ 4 Experimental Setup
We use debugging datasets based on popular LLM coding benchmarks, HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) along with their extended evaluation suites with more unit tests as proposed by Liu et al. (2024b). We describe their construction below, with further details in Sec. A.1.
HE+Fix. We take HumanEvalFix (Muennighoff et al., 2024), which has human-introduced errors, and following Ni et al. (2024), augment each problem’s test suite with overlapping problems from EvalPlus Liu et al. (2024b). This yields 158 problems, each with a faulty solution and an expanded suite of private UTs for evaluation – i.e., UTs not used for debugging.
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark Austin et al. (2021); Liu et al. (2024b); Ni et al. (2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Sec. A.1). Then, we sample one faulty solution for each problem, resulting in 325 problems with incorrect solutions, representing realistic errors made by LLMs. We find such errors are generally harder for models to debug (cf. Sec. 5).
MBPP+Fix (Hard). Additionally, we create a different split from MBPP+ with more subtle errors, logical flaws and missing corner cases making it harder to debug. To this end, we follow a similar generation and filtering process as above but only retain faulty code that passes between of unit tests. This results in a total of 170 problems with faulty codes.
@@
Attack Rate. We measure the attacking ability of a UT generator by matching how often the output of a gold reference solution for input differs from that of the buggy code .
Output Accuracy. This metric measures how often the output of a generated unit test is consistent with the problem description, i.e., generates the same output as the reference gold code .
@@
We compare UTGen against the following (prompts in Appendix E):
No UT feedback: In this baseline, we use self-generated or self-critique feedback, following prior work Madaan et al. (2024); Shinn et al. (2024); Chen et al. (2023b). Specifically, we prompt the model to generate an explanation about the bugs in the target code in addition to determining if the target code is correct, i.e., does not have any more bugs.
Randomly-sampled UTs: Here, we relax the requirement for the LLM to generate failing unit tests, and prompt the model to jointly generate valid unit tests (inputs and outputs) based on the task description, irrespective of the target code, i.e., we sample .
@@ 5.1 Intrinsic Evaluation of UT Generation
In Sec. 5.1, we benchmark the zero-shot abilities of different models on UT generation, as well as our improved training method (corresponding to RQ2). Here, we observe a clear tradeoff: prompted and random unit test generation each optimizing output accuracy or attack rate at the cost of the other metric. While randomly sampled UTs have relatively higher output accuracy (i.e., the output of the unit test is correct according to the problem description), the random baseline often lacks the ability to generate UTs with inputs that trigger errors (failing UTs). This can be explained by the fact that it is not conditioned on the faulty code . On the other hand, when models are prompted to generate a UT that breaks the faulty code in the prompted UT baseline, they generally achieve a higher attack rate; however, they lag in terms of output accuracy. For instance, in the case of Qwen2.5 7B, switching from random sampling to prompting failing UTs improves the attack rate by 13.53% but decreases the output accuracy by 16.47%, hence leading to comparable “Acc. Attack” scores (13.53% vs 14.71%). Taken together, this suggests the presence of competing desirable traits of UT generators (cf. Sec. 3.1; RQ1): without training, one baseline often generates trivial unit tests for which the output can easily be predicted (high output accuracy, low attack rate), while the other generates challenging unit tests for which it cannot produce the expected answer (low output accuracy, high attack rate). It also indicates that while models can generate challenging unit test inputs that result in code failure, models struggle with reasoning over correct outputs of failing UT inputs.
@@
We propose to break through this trade-off (thereby addressing RQ2, “How can we improve LLMs’ UT generation abilities?”) by training models with UTGen, where we directly supervise models to have both high output accuracy and high attack rate. Ideally, UTGen should lead to models that generate unit tests in the “sweet spot” of difficulty, where they are hard enough to trigger errors but not so hard that their output cannot be predicted. Sec. 5.1 shows that models trained with UTGen consistently rank highest at jointly generating failing UT inputs with correct outputs (as measured via “Acc. Attack” in Sec. 5.1). For instance, on the strongest coding model in our setting, Qwen2.5 32B, UTGen obtains the highest attack rate, and output accuracy as well as improves Acc. Attack score by 5.3% (absolute) over failing UTs generated by prompting and by 23.34% over randomly generated UTs. Similarly, Llama3.1 improves (in terms of attack rate) over randomly-sampled UTs by 19.61% and the output accuracy of prompted failing UTs by 20.89%, ultimately improving the joint “Acc. Attack” score by up to 7.65%. We report the intrinsic metrics for HE+Fix and MBPP+Fix in Appendix D.
@@
5.2 Generated UTs for Code Debugging
UTs from UTGen are Best for UTDebug.
Addressing RQ3, we measure how effective each type of UT generator is in a downstream debugging evaluation. From Sec. 5.1, we observe that across different LLM families, multi-turn debugging with UTDebug is more effective when using generated UTs from UTGen than debugging with UTs generated by the baselines, as well as debugging without feedback. For instance, with Qwen2.5 32B, after 3 rounds of debugging on MBPP+Fix, UTGen improves over debugging with randomly-sampled UTs by 4.61%, debugging by failing UTs by 3.38%, and debugging without UT feedback by 15.07%. Moreover, on the more challenging MBPP+Fix (Hard) split, UTGen improves over randomly-sampled UTs by 22.35% and the baseline without UT feedback by 12.35%. Given that we use the same underlying LLM and similar feedback templates, the results in Sec. 5.1 show that UTs generated by UTGen provide the most useful and effective feedback. Lastly, we find that LLM debugging without any UT feedback is least effective across model families, thus, establishing that despite noise and issues discussed in Sec. 3.3 and Fig. 3, UTDebug with UT-based feedback is more effective than self-generated feedback. This aligns with past work indicating that LLMs are poor at self-critique Huang et al. (2024); Stechly et al. (2024); Chen et al. (2024).
@@ 5.3 Generated UTs for Code Generation
Setup.
We further demonstrate the effectiveness of generated UTs at judging code correctness via Best-of- sampling (). Given a problem , we sample different code solutions, and for each solution, we generate 3 UTs (via UTGen and baselines in Sec. 4 with test-time scaling). In the end, we collate all generated UTs from each method (removing any duplicates) and choose the generated code solution that passes the most generated UTs. In other words, we use UTGen to rerank 10 generated solutions. In addition to comparing to the random and prompt UT generation baselines, we also rerank according to a state-of-the-art 8B scale reward model Liu et al. (2024a) on HE+ and MBPP+ (two standard code-generation datasets).
Results.
Based on results in Table 3, we find that:
@@ Results.
UTGen is a Better Judge for Code. Across models and datasets, using UTs from UTGen to select code yields the highest performance, outperforming external RM by 4.36%, randomly generated UTs by 3.73%, and prompted UTs by 1.83% on HE+ with Llama3 8B (with similar trends on MBPP+ and using Qwen2.5 7B).
5.4 On Test-time Scaling and Backtracking
| UT Method | Acc. | |
| Randomly-sampled (Qwen2.5 7B) | 34.77 | |
| - Output Test-time Scaling | 30.77 | - 4.0 |
| - Backtracking | 32.61 | - 2.2 |
| UTGen (Qwen2.5 7B) | 37.54 | |
| - Output Test-time Scaling | 26.15 | - 11.4 |
| - Backtracking | 34.38 | - 3.2 |
We show that test-time scaling and backtracking are crucial for LLM debugging with generated UTs. From Table 4, we observe that irrespective of the underlying method for UT generation (either randomly sampled or from UTGen), removing either backtracking or test-time scaling hurts downstream performance. First, we find that removing test-time scaling for predicting the output of UTs decreases the performance of randomly-sampled UTs by 4% and that of UTGen by 11.4%. Note that a larger drop in UTGen’s performance when removing test-time scaling of UT output prediction is consistent with the findings in Sec. 5.1 that models struggle with reasoning over correct outputs for failing UT inputs (more often generated by UTGen). Therefore, test-time scaling in output prediction provides a greater boost for reasoning over challenging inputs Wang et al. (2022), and consequently, removing it yields larger drops for UTGen. Moreover, Table 4 demonstrates that, without validation on other generated unit tests, LLMs tend to overfit on the unit test contained in the feedback, resulting in up to 3.2% drop in performance of UTGen without backtracking.
5.5 Scaling with Number of Generated UTs
5.6 Comparison to Frontier LLMs
@@
Method Attack Rate Out Acc. Acc. Attack GPT-4o Random 23.92 57.25 17.25 Prompted 54.51 60.39 33.33 DeepSeek V3 Random 24.31 63.53 17.84 Prompted 54.31 58.04 33.33 Qwen2.5 32B Random 25.29 47.25 11.37 Prompted 51.96 48.43 29.41 UTGen 56.08 59.22 34.71
In Sec. 5.1 we evaluated open-source LLMs that we trained using UTGen, and in Sec. 3.3 we applied these models in our UTDebug framework for debugging, finding the trained models to be effective. Sec. 5.1 compares stronger frontier models – GPT-4o Achiam et al. (2023) and DeepSeek-V3 Guo et al. (2024b) – to the best open-source model tested (Qwen2.5 32B). Here, we see that even these stronger and much larger frontier models struggle at the task of generating failing unit tests. First, we find that when prompting models with faulty codes, the attack rate of the generated UTs is slightly over 50% on MBPP+Fix (Hard) dataset, showcasing the inherently challenging nature of isolating corner cases and subtle errors in partially correct code. Furthermore, UTGen beats both GPT-4o and DeepSeek-V3 in terms of Acc. Attack, with a gain of over both, as well as in terms of Attack Rate. On output accuracy, UTGen is better than DeepSeek-V3 and slighlty worse than GPT-4o. Overall, we find that using UT-based feedback even with frontier LLMs only triggers and identifies errors 1 in 3 times, leaving tremendous room for growth in the ability of models to identify partially incorrect code.
@@
6 Discussion and Conclusion
Our work on UTGen contributes to the broader landscape of verification and feedback generation for LLM-generated code. While recent work has focused on training verifiers to provide feedback Mahan et al. (2024); Zhang et al. (2024), a key challenge remains in obtaining high-quality feedback signals for debugging. UTGen addresses this by directly generating unit tests that can identify problems in code, complementing existing work on how to effectively incorporate and present feedback for debugging Chen et al. (2023b); Zhong et al. (2024) along with test-time scaling and backtracking incorporated in UTDebug. Our results demonstrate that without a quality signal to determine code correctness and/or how a faulty code is failing (in the form of unit tests), using LLMs to generate feedback and debug still proves to be challenging. This is one of the first efforts in this direction, and we hope to spark more interest in future work toward LLM-generated unit tests (both input and outputs) that reveal the model’s coding errors.
Our approach connects to and complements recent work on handling real-world programming issues. While approaches designed for SWEBench Jimenez et al. (2024) focus on fixing known issues from GitHub by understanding and implementing fixes for bug reports, UTGen addresses the upstream challenge of automatically discovering potential issues in new code through test generation. Both tasks share a core challenge: determining the expected behavior of code without access to correct implementations. This connects to the fundamental concept of simulatability from computability theory Rice (1953), where we ask whether a program can predict the behavior of another program. Recent work such as Jain et al. (2024) shows that while LLMs can often simulate existing code by tracing execution steps, they struggle more with predicting correct outputs from specifications alone. Our results align with these findings – while UTGen can generate test inputs that trigger errors (high attack rate), predicting correct expected outputs remains challenging (lower output accuracy). This suggests that improving LLMs’ ability to reason about intended program behavior from specifications remains a crucial direction for future work. Nevertheless, we find that the modifications made to debugging in UTDebug help boost UTGen’s accuracy and account for noise, leading to downstream gains.
Conclusion.
@@
We first identified a key trade-off between attack rate and output prediction accuracy when predicting unit tests with models. In light of this trade-off, we introduced UTGen, a new method for creating training data and teaching models to produce unit tests. This allows us to train models to produce better unit tests, as measured by intrinsic metrics like attack rate and output accuracy. Moreover, finding that existing data contains large numbers of easy errors, we introduce a new subset of data with challenging and hard-to-diagnose errors. To enable debugging with automated unit tests, we propose UTDebug, wherein we augment our predictor’s accuracy with test-time scaling and regularize it using a cross-validation and back-tracking procedure that prevents it from overfitting to a narrow or incorrect unit test. This, combined with UTGen, results in consistent increases in debugging performance across models, and can be used to improve code generation via best-of-N sampling.

### v3

@@ 3.3 UTDebug: Generated Unit Tests for Debugging
Building on past work that has shown the benefits of scaling up inference-time compute (Wang et al., 2022; Lightman et al., 2023; Snell et al., 2024), we improve UT output accuracy by allocating additional computation to the problem. Specifically, we use self-consistency (SC; Wang et al., 2022), whereby, for a given UT input, we sample output completions (including CoT rationales) and take the most common final UT output (majority vote) as the final answer, as shown in Fig. 3 (top-right). To further boost output accuracy, consistent with Prasad et al. (2024), we upsample UT inputs and only retain those where the final answer gets over 50% of the votes (i.e., 4 votes), discarding the unit test otherwise.
@@ 4 Experimental Setup
We demonstrate the effectiveness of UTGen on three 7-8B scale LLMs across different model families that are adept at coding, namely, Llama3 8B Instruct (AI@Meta, 2024), Llama3.1 8B Instruct (Dubey et al., 2024), Qwen 2.5 Coder 7B Instruct in addition to the stronger Qwen 2.5 Coder 32B Instruct model (Hui et al., 2024).
@@
We use debugging datasets based on popular LLM coding benchmarks, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) along with their extended evaluation suites with more unit tests as proposed by (Liu et al., 2024b). We describe their construction below, with further details in Sec. A.1.
HE+Fix. We take HumanEvalFix (Muennighoff et al., 2024), which has human-introduced errors, and following Ni et al. (2024), augment each problem’s test suite with overlapping problems from EvalPlus (Liu et al., 2024b). This yields 158 problems, each with a faulty solution and a suite of private UTs for evaluation – i.e., UTs not used for debugging.
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark (Austin et al., 2021; Liu et al., 2024b; Ni et al., 2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Sec. A.1). Then, we sample one faulty solution for each problem, resulting in 325 problems with realistic incorrect solutions generated by LLMs. We find such errors are generally harder for models to debug (cf. Sec. 5).
MBPP+Fix (Hard). Additionally, we create a different split from MBPP+ with more subtle errors, logical flaws and missing corner cases making it harder to debug. To this end, we follow a similar generation and filtering process as above but only retain faulty code that passes between of unit tests, resulting in 170 problems with faulty codes.
@@
Attack Rate. We measure the attacking ability of a UT generator by the frequency with which the output of a gold solution for input differs from that of the buggy code .
Output Accuracy. Measures how often the output of a generated unit test is consistent with the problem description, i.e., generates the same output as the reference code .
@@
We compare UTGen against the following (prompts in Appendix H):
No UT feedback: In this baseline, we use self-generated or self-critique feedback, following prior work (Madaan et al., 2024; Shinn et al., 2024; Chen et al., 2023b). Specifically, we prompt the model to generate an explanation about the bugs in the target code in addition to determining if the target code is correct, i.e., does not have any more bugs.
Randomly-sampled UTs: Consistent with Chen et al. (2023a), we relax the requirement for the LLM to generate failing unit tests, and prompt the model to jointly generate valid unit tests (inputs and outputs) based on the task description, irrespective of the target code, i.e., we sample . We further expand on the comparison with Chen et al. (2023a) in Appendix E.
@@
Note that all UT-generation baselines along with UTGen utilize the test-time scaling and backtracking approaches of UTDebug highlighted in Sec. 3.3 and exhibit similar run times.
@@ 5.1 Intrinsic Evaluation of UT Generation
In Sec. 5.1, we benchmark the zero-shot abilities of different models on UT generation, as well as our improved training method (corresponding to RQ2).
Here, we observe a clear tradeoff: prompted and random unit test generation each optimizing output accuracy or attack rate at the cost of the other metric.
While randomly sampled UTs have relatively higher output accuracy (i.e., the output of the unit test is correct according to the problem description), the random baseline often lacks the ability to generate UTs with inputs that trigger errors (failing UTs).
This can be explained by the fact that it is not conditioned on the faulty code .
On the other hand, when models are prompted to generate a UT that breaks the faulty code in the prompted UT baseline, they generally achieve a higher attack rate; however, they lag in terms of output accuracy.
For instance, in the case of Qwen2.5 7B, switching from random sampling to prompting failing UTs improves the attack rate by 13.53% but decreases the output accuracy by 16.47%, hence leading to comparable Acc. Attack scores (13.53% vs 14.71%).
Taken together, this suggests the presence of competing desirable traits of UT generators (cf. Sec. 3.1; RQ1): without training, one baseline often generates trivial unit tests for which the output can easily be predicted (high output accuracy, low attack rate), while the other generates challenging unit tests for which it cannot produce the expected answer (low output accuracy, high attack rate).
It also indicates that while models can generate challenging unit test inputs that result in code failure, models struggle with reasoning over correct outputs of failing UT inputs.
@@
We propose to break through this trade-off (thereby addressing RQ2, “How can we improve LLMs’ UT generation abilities?”) by training models with UTGen, where we directly supervise models to have both high output accuracy and high attack rate. Ideally, UTGen should lead to models that generate unit tests in the “sweet spot” of difficulty, where they are hard enough to trigger errors but not so hard that their output cannot be predicted. Sec. 5.1 shows that models trained with UTGen consistently rank highest at jointly generating failing UT inputs with correct outputs (as measured via Acc. Attack in Sec. 5.1). For instance, on the strongest coding model in our setting, Qwen2.5 32B, UTGen obtains the highest attack rate, and output accuracy as well as improves Acc. Attack score by 5.3% (absolute) over failing UTs generated by prompting and by 23.34% over randomly generated UTs. Similarly, Llama3.1 improves (in terms of attack rate) over randomly-sampled UTs by 19.61% and the output accuracy of prompted failing UTs by 20.89%, ultimately improving the joint Acc. Attack score by up to 7.65%. We report the intrinsic metrics for HE+Fix and MBPP+Fix in Appendix D.
5.2 Generated UTs for Code Debugging
UTs from UTGen are Best for UTDebug.
Addressing RQ3, we measure how effective each type of UT generator is in a downstream debugging evaluation. From Sec. 5.2,
@@ 5.2 Generated UTs for Code Debugging
we observe that across different LLM families, multi-turn debugging with UTDebug is more effective when using generated UTs from UTGen than debugging with UTs generated by the baselines, as well as debugging without feedback. For instance, with Qwen2.5 32B, after 3 rounds of debugging on MBPP+Fix, UTGen improves over debugging with randomly-sampled UTs by 4.61%, debugging by failing UTs by 3.38%, and debugging without UT feedback by 15.07%. Moreover, on the more challenging MBPP+Fix (Hard) split, UTGen improves over randomly-sampled UTs by 22.35% and the baseline without UT feedback by 12.35%. Given that we use the same underlying LLM and similar feedback templates, the results in Sec. 5.2 show that UTs generated by UTGen provide the most useful and effective feedback. Lastly, we find that LLM debugging without any UT feedback is least effective across model families, thus, establishing that despite noise and issues discussed in Sec. 3.3 and Fig. 3, UTDebug with UT-based feedback is more effective than self-generated feedback. This aligns with past work indicating that LLMs are poor at self-critiquing (Huang et al., 2024).
@@ 5.3 Generated UTs for Code Generation
We further demonstrate the effectiveness of generated UTs at judging code correctness via Best-of- sampling (). Given a problem , we sample different code solutions, and for each solution, we generate 3 UTs (via UTGen and baselines in Sec. 4 with test-time scaling). In the end, we collate all generated UTs from each method (removing any duplicates) and choose the generated code solution that passes the most generated UTs. In other words, we use UTGen to rerank 10 generated solutions. In addition to comparing to the random and prompt UT generation baselines, we also rerank according to a state-of-the-art 8B scale reward model (Liu et al., 2024a) on HE+ and MBPP+ (two standard code-generation datasets). Results in Table 3 reveal that:
@@
UTGen is a Better Judge for Code. Across models and datasets, using UTs from UTGen to select code yields the highest performance, outperforming external RM by 4.36%, randomly generated UTs as employed by Chen et al. (2023a) by 3.73%, and prompted UTs by 1.83% on HE+ with Llama3 8B (with similar trends on MBPP+ and using Qwen2.5 7B).
5.4 Comparison to Unit Test Generation by Frontier LLMs
@@
Method Attack Rate Out Acc. Acc. Attack GPT-4o Random 23.92 57.25 17.25 Prompted 54.51 60.39 33.33 DeepSee V3 Random 24.31 63.53 17.84 Prompted 54.31 58.04 33.33 Qwen2. 32B Random 25.29 47.25 11.37 Prompted 51.96 48.43 29.41 UTGen 56.08 59.22 34.71
In Sec. 5.1 we evaluated open-source LLMs that we trained using UTGen, and in Sec. 3.3 we applied these models in our UTDebug framework for debugging, finding the trained models to be effective. Sec. 5.2 compares stronger frontier models – GPT-4o (Hurst et al., 2024) and DeepSeek-V3 (Guo et al., 2024b) – to the best open-source model tested (Qwen2.5 32B). Here, we see that even these stronger and much larger frontier models struggle at the task of generating failing unit tests. First, we find that when prompting models with faulty codes, the attack rate of the generated UTs is slightly over 50% on MBPP+Fix (Hard) dataset, showcasing the inherently challenging nature of isolating corner cases and subtle errors in partially correct code. Furthermore, UTGen beats both GPT-4o and DeepSeek-V3 in terms of Acc. Attack, with a gain of over both, as well as in terms of Attack Rate. On output accuracy, UTGen is better than DeepSeek-V3 and slightly worse than GPT-4o. Overall, we find that using UT-based feedback even with frontier LLMs only triggers and identifies errors 1 in 3 times, leaving tremendous room for growth in the ability of models to identify partially incorrect code.
@@
5.5 Importance of Unit Tests with a Frontier Debugger
| Debugger | UT Method (Model) | MBPP+Fix (Hard) |
| GPT4o | No UT (GPT-4o) | 34.71 |
| No UT (Qwen 32B) | 32.94 | |
| Prompted (Qwen 32B) | 45.88 | |
| UTGen (Qwen 32B) | 59.69 |
Finally, we study the importance of unit tests when using a more powerful, frontier model for debugging. To this end, we use GPT-4o (Hurst et al., 2024) as the debugger and evaluate its performance on the MBPP+Fix (Hard) benchmark. We compare a baseline that does not use UT-execution feedback, where GPT-4o generates its own feedback, or uses feedback generated by Qwen2.5 32B (without any unit tests) against scenarios where it uses UT feedback from the smaller Qwen2.5 32B model, both from a prompted baseline and from our proposed UTGen. The results, presented in Table 5, show that even with a significantly stronger debugger model, feedback from high-quality unit tests remains crucial for effective debugging – both the prompted baseline and UTGen outperform the “no UT” baselines by 11.17% and 25% (absolute), respectively. This demonstrates not only that UT feedback is vital even for frontier models, but also that the quality of the unit tests is a key factor, as UTGen outperforms the prompted baseline by a large margin of 13.8%.
6 Conclusion
@@
We first identified a key trade-off between attack rate and output prediction accuracy when predicting unit tests with models. In light of this trade-off, we introduced UTGen, a new method for creating training data and teaching models to produce unit tests. This allows us to train models to produce better unit tests, as measured by intrinsic metrics like attack rate and output accuracy. Moreover, finding that existing data contains large numbers of easy errors, we introduce a new subset of data with challenging and hard-to-diagnose errors. To enable debugging with automated unit tests, we propose UTDebug, wherein we augment our predictor’s accuracy with test-time scaling and regularize it using a cross-validation and back-tracking procedure that prevents it from overfitting to a narrow or incorrect unit test. Additionally, in Appendix F, we demonstrate that these gains persist as we scale the number of generated UTs across datasets. This, combined with UTGen, results in consistent increases in debugging performance across models, and can be used to improve code generation via best-of-N sampling. Finally, we note that UTGen is complementary to work on handling real-world programming issues such as Jimenez et al. (2024), which focus on debugging GitHub issues raised by people; in the future, such issues could be identified by a unit test generator such as UTGen.